a constant power p is applied on a body of mass m by a vairiable force in same direction of displacement find relation 
1=Velocity respect to time?
2=Displacement with respect to time?
please explain in detail

Grade:11

1 Answers

Mayank
4 years ago
  1. ∫p=W wrt to time
W- 1/2mv2
pt=1/2mv2
  1. v=dx/dt
pt=1/2,m(dx/dt)2
if x is given in for of equation first find derivative then square
else integrat and solve
 
​​